how do you find five consecutive odd integers such that the sum of the first and the fifth is one less than three times the fourth
Answer by tommyt3rd(5050)

consecutive integer will be these...
1:n
2:n+2
3:n+4
4:n+6
5:n+8
n + (n+8) = 3(n+6)-1
2n+8=3n+18-1
-n=17-8
n=-9
they are...
-9, -7, -5, -3, -1